Write an equation for a sine curve that has the given amplitude, period, and phase shift.
2, pi/3, pi/3

Answer by lwsshak3(11628)

Write an equation for a sine curve that has the given amplitude, period, and phase shift.
2, pi/3, pi/3
**
Equation for sin function:
y=Asin(Bx-C), A=amplitude, period=2π/B, C= Phase shift
For given problem:
A=2
period=π/3
π/3=2π/B
B=6
C=π/3
Given equation for sin function:
y=2sin(6x-π/3)